A city bus arrives at your bus stop every 8 minutes and follows a uniform distribution. The average wait time is

1 minute.

2 minutes.

3 minutes.

4 minutes.

5 minutes.

None of the above.
Answer #1

The distribution is uniform.

The wait time after bus leaves = 8 mins.

The wait time after one min = 7 mins.

The wait time after two mins = 6 mins.

......

The wait time after 7 mins = 1 mins.

The function is f(t) = 8 - t and the summation is

gif.latex?%5Cint_%7B0%7D%5E%7B8%7D%20%28

= 8t - t2/2 |08

= 8*8 - 82/2 - 0 + 0

= 64 - 32

= 32.

=> Average wait time = 32/8 = 4 minutes.